OpenStudy (anonymous):

In the Pythagorean Theorem, the “c” represents the difference of the legs of the right triangle. True/False

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (midhun.madhu1987):

figure?

OpenStudy (anonymous):

There is no figure.

OpenStudy (skullpatrol):

$$\Huge a^2+b^2=c^2$$

OpenStudy (anonymous):

So, it's true?

OpenStudy (midhun.madhu1987):

no

OpenStudy (skullpatrol):

c is the sum of the squares, right?

OpenStudy (anonymous):

So, the c represents the hypotenuse?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

c is a letter it can represent anything

OpenStudy (anonymous):

sometimes it is the speed of light as in \(E=MC^2\)

OpenStudy (skullpatrol):

$$\Huge a^2+b^2=c^2$$ This is the standard abc representation of the theorem

OpenStudy (anonymous):

@skullpatrol knows better \[\Huge a^2+b^2=c^2\] is not a theorem, it is an equation it can either be true, for example if \(a=3,b=4,c=5\) or it can be false, for example \(a=10,b=1,c=\pi\)

OpenStudy (anonymous):

a theorem is not an equation a theorem says "if something is true, then something else is true" the "something else" part can be an equation

OpenStudy (skullpatrol):

theorem: A statement that is shown to be true by a logically developed argument.

OpenStudy (skullpatrol):



OpenStudy (anonymous):

if you label the sides of a right triangle, then fine, that is the "if" part

OpenStudy (anonymous):

"if \(a, b, c\) are the lengths of the legs resp hypotenuse of a right triangle, then \(a^2+b^2=c^2\)" is a theorem \[a^2+b^2=c^2\] is not

OpenStudy (skullpatrol):

The question asks: In the Pythagorean Theorem, the “c” represents the difference of the legs of the right triangle. True/False You will have to make some standard assumptions to make any sense of this question, if you don't you cannot answer it.

OpenStudy (anonymous):

frankly the math teacher who wrote this nonsense should be removed from the classroom no wonder students don't know the difference between a theorem and an equation, let alone understand what a variable is

OpenStudy (skullpatrol):

True dat^

OpenStudy (skullpatrol):

But until that happens we have to work with what we got :-)

OpenStudy (anonymous):

yeah true dat too (unfortunately)